# Leveraging Pre-Trained Visual Transformers for Multi-Band Photometric Light Curve Classification

### v1

@@ 1 Introduction
Machine learning and deep learning techniques have played a central role in advancing the classification of light curves (Pichara et al. 2012; Kim et al. 2011; Villar et al. 2019; Becker et al. 2020; Cabrera-Vives et al. 2024). Traditional machine learning approaches have been highly effective in classifying variable stars and transient events by relying on hand-engineered features extracted from light curves. For example, Richards et al. (2011) and Kim et al. (2014) employed statistical and time-domain features to enhance classification accuracy. Karpenka et al. (2013) combined parametric functional fitting of supernova light curves with a machine learning algorithm, yielding accurate and robust classifications. Lochner et al. (2016) developed a multi-stage pipeline that evaluates feature extraction techniques and machine learning models. Boone (2019) improved classification by training a boosted decision tree on features extracted from the augmented light curves, achieving state-of-the-art performance in the LSST Astronomical Time-Series Classification Challenge (PLAsTiCC; Kessler et al. 2019). However, this feature engineering process is resource-intensive and requires domain knowledge to be applied (Donalek et al. 2013; Graham et al. 2014; Nun et al. 2017; Sánchez-Sáez et al. 2021). Deep learning addresses this limitation by automatically learning relevant representations directly from light curves, removing the need for manual feature extraction. This advantage has led to the widespread adoption of recurrent neural networks (RNNs; Rumelhart et al. 1986), long short-term memory networks (LSTMs; Hochreiter 1997), and gated recurrent units (GRUs; Cho et al. 2014) (Charnock & Moss 2017; Naul et al. 2018; Muthukrishna et al. 2019; Carrasco-Davis et al. 2019; Möller & de Boissière 2020; Becker et al. 2020; Jamal & Bloom 2020; Gómez et al. 2020; Donoso-Oliva et al. 2021), as well as transformer-based models, which have shown great promise in handling astronomical time series data (Morvan et al. 2022; Pan et al. 2022; Pimentel et al. 2022; Donoso-Oliva et al. 2023; Moreno-Cartagena et al. 2023; Cádiz-Leyton et al. 2024a, b; Allam Jr & McEwen 2024; Cabrera-Vives et al. 2024), making them increasingly popular for astronomical classification tasks.
@@
Foundational models, which are trained only once on massive datasets in a self-supervised manner (pre-training) to learn general domain representations and later adapted for specific tasks (fine-tuning), have demonstrated exceptional generalization across diverse downstream applications in natural language processing and computer vision, such as text generation (Brown et al. 2020), machine translation (Vaswani et al. 2017), image classification (Dosovitskiy et al. 2021), and object detection (Carion et al. 2020), among others. Notable examples include GPT (Radford 2018) and BERT (Devlin et al. 2019) for text, as well as Vision Transformer (ViT; Dosovitskiy et al. 2021) for images, each specializing in a single modality. In the context of astronomical time series data, Astromer (Donoso-Oliva et al. 2023) represents an early effort to pre-train a transformer on large-scale real-world light curve data, positioning it as a potential foundational model for time-domain astronomy. However, such pre-training requires significant computational resources, optimized architectural design, and carefully curated datasets, often taking several days to converge.
@@
In this work, we present the use of a pre-trained VT model called Swin Transformer V2 (SwinV2; Liu et al. 2022), which we evaluate on two datasets: the public Catalog of Variable Stars (Alcock et al. 2003), derived from the Massive Compact Halo Object (MACHO; Alcock et al. 2000) survey, and the first round of the recent Extended LSST Astronomical Time-Series Classification Challenge (ELAsTiCC [^1] The DESC ELAsTiCC Challenge.; Narayan & Team 2023), a multi-band dataset of photometric light curves for variable stars and supernovae designed to simulate the observations of the future LSST. Our work builds upon the methodology proposed by Li et al. (2024), which focuses on benchmark datasets for time series. We applied this method to light curves, performed hyperparameter tuning, and evaluated its performance using both single- and multi-band configurations for the MACHO dataset and six bands for the ELAsTiCC dataset. Additionally, we assess its performance in scenarios with a limited number of astronomical objects per class and imbalanced datasets. To our knowledge, this is the first study to apply a pre-trained VT to raw multi-band light curves in the time domain. The key contributions of this work are as follows:
  • Paving the way for the use of pre-trained VT models to analyze and classify light curves.
  • Achieving competitive results compared to state-of-the-art methods using only light curves.
  • Providing a robust solution capable of leveraging multiple bands simultaneously.
  • Presenting a simple and effective approach for the classification of photometric light curves.
@@ 2.1 Multi-band light curves to images
To ensure that all light curve images are on the same scale and maintain consistent x- and y-limits across different images, we apply Min-Max normalization to each entire light curve (considering all bands). This approach preserves the relative differences among bands within each light curve while scaling time and flux values across all light curves to the range [00, $1$]. Additionally, the flux errors are scaled using the same Min-Max factor applied to the flux, maintaining the relative uncertainty information. This method ensures that all images are visually comparable, with uniform axis limits and scale throughout the dataset.
@@ 2.2 Visual transformer models
SwinV2 is a hierarchical VT architecture designed for efficient and scalable image processing (Liu et al. 2022). Unlike ViT (Dosovitskiy et al. 2021), which partitions an image into non-overlapping segments called patches, applying global self-attention to all of them simultaneously, the first Swin Transformer paper (Liu et al. 2021) introduces two key innovations that allow it to manage both local and global self-attention: shifted windows and hierarchical feature representation. These features enable the model to handle high-resolution images with greater computational efficiency. Furthermore, SwinV2 incorporates scaled cosine attention and logarithmic continuous position bias, improving both training stability and performance in tasks such as image classification, object detection, semantic segmentation, and video action classification. We selected this model as the foundation of our approach because its hierarchical structure is particularly effective at capturing detailed information at the pixel level, which is essential for interpreting points and lines within the light curve image. Figure 2 illustrates the model’s architecture and hierarchical processing. In the following items, we break down its components and functionality in detail.
@@
  1. 1.
    Patch Splitting and Tokenization: Similar to ViT, Swin first partitions the input image into non-overlapping patches. However, unlike ViT, Swin progressively reduces the patch size as the image propagates through the Swin Transformer block (explained in item 6). Given an input image of $256{\times}256$ pixels with 3 color channels (e.g., RGB) and a patch size of $4{\times}4$, the patch partitioning process yields $64{\times}64$ patches. This is equivalent to applying a convolutional kernel of size $4{\times}4$ with a stride of $4{\times}4$ over the image. Each patch contains $4{\times}4{\times}3{=}48$ values, which represent the RGB pixel intensities within that region. These values are then flattened and projected into a higher-dimensional space via a linear transformation, producing a token map of dimensions $64{\times}64{\times}C$, where $C$ is the embedding dimension. In this context, a token refers to the vector representation of a patch after the linear transformation. This process is functionally analogous to applying a two-dimensional convolution layer (Conv2d). The resulting token map is subsequently processed through the Swin Transformer block for hierarchical processing.
  2. 2.
    Window Multi-head Self-Attention (W-MSA): The first step of the Swin Transformer block partitions the tokens into non-overlapping windows of size $M{\times}M$. Unlike traditional Multi-head Self-Attention (MSA), which computes attention globally and has quadratic complexity, W-MSA restricts attention to local windows. This reduces the computational complexity from $\mathcal{O}(N^{2})$ in global attention to $\mathcal{O}(M^{2}N)$ in W-MSA, where $N$ denotes the total number of tokens in the image. Since $M$ is fixed, W-MSA maintains linear complexity with respect to $N$, significantly reducing computational cost while ensuring that each window effectively captures local dependencies, making the approach scalable for high-resolution images.
  3. 3.
    Shifted Window Multi-head Self-Attention (SW-MSA): To address the limited cross-window interactions in W-MSA, SW-MSA introduces a shifted window mechanism. This shifts the windows from the W-MSA step, allowing tokens at the edges of one window to be included in the centers of overlapping windows. This process enables information exchange between neighboring windows. By alternating W-MSA and SW-MSA within each Swin Transformer block, the model captures global context while preserving the computational advantages of local attention. For a more detailed explanation of the W-MSA and SW-MSA mechanisms, please refer to the first Swin Transformer paper (Liu et al. 2021).
  4. 4.
    Post Normalization and Scaled Cosine Attention: Most VT models, including the original Swin Transformer, apply layer normalization before the attention mechanism. However, as model depth increases, deeper layers develop larger activation magnitudes due to residual connections, leading to training instability. To address this, SwinV2 adopts residual post-normalization, where normalization is applied after the attention mechanism and the output of the Swin Transformer block, preventing the accumulation of large activations. Additionally, traditional dot-product attention in large models can cause certain heads to focus excessively on a few pixel pairs, particularly in the residual post-normalization setting. To mitigate this, SwinV2 introduces scaled cosine attention, which replaces dot-product similarity with cosine similarity, normalizing attention scores and improving training stability.
  5. 5.
    Log-spaced Continuous Position Bias: SwinV2 also replaces the relative position bias with a log-spaced continuous position bias. This design prioritizes local token interactions while gradually reducing the influence of distant tokens, enabling the model to generalize better across different image resolutions.
  6. 6.
    Hierarchical Representation with Patch Merging: Following the operations described in items 2 to 5, which constitute the Swin Transformer block, the model applies patch merging to create a hierarchical representation. This operation reduces the spatial resolution of the token map while increasing its embedding dimension. For instance, a token map of size $64{\times}64{\times}96$ is downsampled to $32{\times}32{\times}192$, where the resolution is halved, and the embedding dimension is doubled. By incorporating this output into subsequent Swin Transformer blocks and further applying patch merging, the hierarchical architecture enables the model to effectively capture multi-scale features.
  7. 7.
    Downstream Task: After the hierarchical stages, the final token representations undergo adaptive average pooling, which computes the mean across all tokens of size $\frac{H}{32}\times\frac{W}{32}$, preserving their aggregated embedding information ($8C$). This process allows each image to be represented as a vector of dimension $8C$, which is then passed through a fully connected linear layer for classification or other downstream tasks.
@@ 3.1.1 MACHO labeled
The public Catalog of Variable Stars [^2] MACHO Labeled Data of the Variable Stars. (Alcock et al. 2003) contains labeled light curves from the MACHO survey, which observed variable stars in the Large Magellanic Cloud (LMC) and the Galactic Bulge over multiple years, capturing their photometric variability in the B and R bands. This dataset has a median cadence of $0.75$ days with a standard deviation of $5.03$ days, considering both bands. The labeled portion we use corresponds to the same one used in Donoso-Oliva et al. (2023), which was categorized into six distinct classes, comprising 20 8942089420\,89420 894 variable stars. We will refer to this as the Full dataset.
Following the approach in Donoso-Oliva et al. (2023), we use scenarios with few samples per class (spc) to train the model and assess its ability to learn from limited labeled data while keeping the test set unchanged. From the Full dataset, we extracted 20 spc and 500 spc as our new training datasets. As in the referenced study, the IDs for each fold were selected randomly from the Full dataset. For example, the IDs for 20 and 500 spc in the first fold differ from those in the second fold. Our training, validation, and test sets in each fold use the same IDs that were used in the referenced study. The test set contains $100$ light curves per class and remains fixed across all scenarios (20 spc, 500 spc, and Full dataset). The training dataset is divided into three stratified folds, maintaining an 80/20 split for training and validation. This results in $96$ and $24$ light curves for the 20 spc scenario, 2 40024002\,4002 400 and $600$ for the 500 spc scenario, and 16 2321623216\,23216 232 and 4 06240624\,0624 062 for the Full dataset. Figure 3 illustrates the class distribution in the training and validation sets for the Full dataset.
@@ 3.1.2 ELAsTiCC Challenge
The second dataset used in this study comes from the ELAsTiCC [^3] ELAsTiCC Challenge Data Portal. challenge (Narayan & Team 2023), designed to simulate the observations of the future LSST at the Vera C. Rubin Observatory. It contains 1 845 14618451461\,845\,1461 845 146 simulated light curves, each represented across six bands: u, g, r, i, z, and Y, covering the visible and near-infrared spectrum. The data includes multiple observations of each object spread over several simulated years. ELAsTiCC spans a diverse array of transient and variable stellar events, categorized into $32$ distinct classes, with a median cadence of $0.93$ days and a standard deviation of $16.18$ days, considering all six bands. Specifically, we used the first ELAsTiCC campaign, which was used in Cabrera-Vives et al. (2024), and adopted the same class distribution as their study, encompassing $20$ different categories. Additionally, similar to the approach outlined in the referenced study, we applied the PHOTFLAG [^4] Description of the data variables from the first round of ELAsTiCC. key to filter out saturated data. For each light curve, we also considered both alerts and forced photometry points, beginning $30$ days prior to the first alert up to the final detection. Non-detections following the last detection were omitted from the dataset.
Our training, validation, and test sets were the same as those used in Cabrera-Vives et al. (2024). The test set contains 1 00010001\,0001 000 light curves from each class, ensuring a balanced representation. The remaining data is divided into five stratified folds, maintaining an 80/20 ratio for training and validation. This results in 1 460 11714601171\,460\,1171 460 117 and 365 029365029365\,029365 029 light curves in the training and validation sets, respectively. Figure 4 shows the distribution of the classes in the training and validation sets.

### v2

@@ 1 Introduction
Machine learning and deep learning techniques have played a central role in advancing the classification of light curves (Pichara et al. 2012; Kim et al. 2011; Villar et al. 2019; Becker et al. 2020; Cabrera-Vives et al. 2024). Traditional machine learning approaches have been highly effective in classifying variable stars and transient events by relying on hand-engineered features extracted from light curves. For example, Richards et al. (2011) and Kim et al. (2014) employed statistical and time-domain features to enhance classification accuracy. Karpenka et al. (2013) combined parametric functional fitting of supernova light curves with a machine learning algorithm, yielding accurate and robust classifications. Lochner et al. (2016) developed a multi-stage pipeline that evaluates feature extraction techniques and machine learning models. Boone (2019) improved classification by training a boosted decision tree on features extracted from the augmented light curves, achieving state-of-the-art performance in the LSST Astronomical Time-Series Classification Challenge (PLAsTiCC; Kessler et al. 2019). However, this feature engineering process is resource-intensive and requires domain knowledge to be applied (Donalek et al. 2013; Graham et al. 2014; Nun et al. 2017; Sánchez-Sáez et al. 2021). Deep learning addresses this limitation by automatically learning relevant representations directly from light curves, removing the need for manual feature extraction. This advantage has led to the widespread adoption of recurrent neural networks (RNNs; Rumelhart et al. 1986), long short-term memory networks (LSTMs; Hochreiter & Schmidhuber 1997), and gated recurrent units (GRUs; Cho et al. 2014) (Charnock & Moss 2017; Naul et al. 2018; Muthukrishna et al. 2019; Carrasco-Davis et al. 2019; Möller & de Boissière 2020; Becker et al. 2020; Jamal & Bloom 2020; Gómez et al. 2020; Donoso-Oliva et al. 2021), as well as transformer-based models, which have shown great promise in handling astronomical time series data (Morvan et al. 2022; Pan et al. 2022; Pimentel et al. 2022; Donoso-Oliva et al. 2023; Moreno-Cartagena et al. 2023; Cádiz-Leyton et al. 2024, 2025; Allam Jr & McEwen 2024; Cabrera-Vives et al. 2024), making them increasingly popular for astronomical classification tasks.
@@
Foundational models, which are trained only once on massive datasets in a self-supervised manner (pre-training) to learn general domain representations and later adapted for specific tasks (fine-tuning), have demonstrated exceptional generalization across diverse downstream applications in natural language processing and computer vision, such as text generation (Brown et al. 2020), machine translation (Vaswani et al. 2017), image classification (Dosovitskiy et al. 2021), and object detection (Carion et al. 2020), among others. Notable examples include GPT (Radford et al. 2018) and BERT (Devlin et al. 2019) for text, as well as Vision Transformer (ViT; Dosovitskiy et al. 2021) for images, each specializing in a single modality. In the context of astronomical time series data, Astromer (Donoso-Oliva et al. 2023) represents an early effort to pre-train a transformer on large-scale real-world light curve data, positioning it as a potential foundational model for time-domain astronomy. However, such pre-training requires significant computational resources, optimized architectural design, and carefully curated datasets, often taking several days to converge.
@@
In this work, we present the use of a pre-trained VT model called Swin Transformer V2 (SwinV2; Liu et al. 2022), which we evaluate on two datasets: the public Catalog of Variable Stars (Alcock et al. 2003), derived from the Massive Compact Halo Object (MACHO; Alcock et al. 2000) survey, and the first round of the recent Extended LSST Astronomical Time-Series Classification Challenge (ELAsTiCC [^1] The DESC ELAsTiCC Challenge.; Narayan & Team 2023), a multi-band dataset of photometric light curves for variable stars and supernovae designed to simulate the observations of the future LSST. Our work builds upon the methodology proposed by Li et al. (2024), which focuses on benchmark datasets for time series. We applied this method to light curves, performed hyperparameter tuning, and evaluated its performance using both single- and multi-band configurations for the MACHO dataset and six bands for the ELAsTiCC dataset. Additionally, we assess its performance in scenarios with a limited number of astronomical objects per class and imbalanced datasets. To our knowledge, this is the first study to apply a pre-trained VT to raw multi-band light curves in the time domain.
This paper is organized as follows: In Section 2, we describe the methodology for converting multi-band light curves into images and present the architecture of the SwinV2 visual transformer model. Section 3 details the datasets and experimental setup. In Section 4, we present the results of our classification experiments on the MACHO and ELAsTiCC datasets, including comparisons with established methods. Finally, Section 5 summarizes the main conclusions of this work and suggests directions for future research.
@@ 2.1 Multi-band light curves to images
To ensure that all light curve images are on the same scale and maintain consistent x- and y-limits across different images, we apply Min-Max normalization to each entire light curve (considering all bands). This approach preserves the relative differences among bands within each light curve while scaling time and flux values across all light curves to the range [0, 11]. As a consequence, the global temporal and brightness context across light curves is removed, which encourages the model to focus on light curve shape to discriminate between astronomical objects. Global normalization was avoided, as it suppressed low-amplitude variability and negatively affected classification performance. Axis labels and ticks are also removed before saving the plots. Additionally, the flux errors are scaled using the same Min-Max factor applied to the flux, maintaining the relative uncertainty information. This method ensures that all images are visually comparable, with uniform axis limits and scale throughout the dataset. Visual examples of these representation approaches applied to randomly selected light curves from the test set are provided in Appendix A, in Fig. 10 and Fig. 11, offering a broader view of the differences between the One-band, Overlay, and Grid methods across both the MACHO and ELAsTiCC datasets.
@@ 2.2 Visual transformer models
SwinV2 is a hierarchical VT architecture designed for efficient and scalable image processing (Liu et al. 2022). Unlike ViT (Dosovitskiy et al. 2021), which partitions an image into non-overlapping segments called patches, applying global self-attention to all of them simultaneously, the first Swin Transformer paper (Liu et al. 2021) introduces two key innovations that allow it to manage both local and global self-attention: shifted windows and hierarchical feature representation. These features enable the model to handle high-resolution images with greater computational efficiency. Furthermore, SwinV2 incorporates scaled cosine attention and logarithmic continuous position bias, improving both training stability and performance in tasks such as image classification, object detection, semantic segmentation, and video action classification. We selected this model as the foundation of our approach because its hierarchical structure is particularly effective at capturing detailed information at the pixel level, which is essential for interpreting points and lines within the light curve image. Figure 2 illustrates the model’s architecture and hierarchical processing. The key stages of this architecture are detailed below.
@@
The first step in the SwinV2 architecture converts the input image into a sequence of tokens through patch splitting and tokenization. This process begins by partitioning the input image into non-overlapping patches, similar to ViT. However, unlike ViT, Swin progressively reduces the patch size as the image propagates through the Swin Transformer block. Given an input image of $256{\times}256$ pixels with 3 color channels (e.g., RGB) and a patch size of $4{\times}4$, the patch partitioning process yields $64{\times}64$ patches. This is equivalent to applying a convolutional kernel of size $4{\times}4$ with a stride of $4{\times}4$. Each patch contains $4{\times}4{\times}3{=}48$ values that represent the RGB pixel intensities within that region. These values are then flattened and projected into a higher-dimensional space via a linear transformation, producing a token map of dimensions $64{\times}64{\times}C$, where $C$ is the embedding dimension. In this context, a token refers to the vector representation of a patch after the linear transformation. This process is functionally analogous to applying a two-dimensional convolution layer (Conv2d). The resulting token map is then processed through the Swin Transformer block for hierarchical processing.
The Window Multi-head Self-Attention (W-MSA) step partitions the tokens into non-overlapping windows of size $M\times M$. Unlike traditional Multi-head Self-Attention (MSA), which computes attention globally with quadratic complexity, W-MSA restricts attention to local windows. This reduces the computational complexity from $\mathcal{O}(N^{2})$ in global attention to $\mathcal{O}(M^{2}N)$ in W-MSA, where $N$ denotes the total number of tokens in the image. Since $M$ is fixed, W-MSA maintains linear complexity with respect to $N$, significantly reducing computational cost while ensuring that each window effectively captures local dependencies, making the approach scalable for high-resolution images.
To address the limited cross-window interactions in W-MSA, the Shifted Window Multi-head Self-Attention (SW-MSA) mechanism shifts the windows from the W-MSA step, allowing tokens at the edges of one window to be included in the centers of overlapping windows. This process enables information exchange between neighboring windows. By alternating W-MSA and SW-MSA within each Swin Transformer block, the model captures global context while preserving the computational advantages of local attention. For a more detailed explanation of the W-MSA and SW-MSA mechanisms, please refer to the first Swin Transformer paper (Liu et al. 2021).
Most VT models, including the original Swin Transformer, apply layer normalization before the attention mechanism. However, as model depth increases, deeper layers develop larger activation magnitudes due to residual connections, leading to training instability. To address this, SwinV2 adopts residual post-normalization, where normalization is applied after the attention mechanism and the output of the Swin Transformer block, preventing the accumulation of large activations. Additionally, traditional dot-product attention in large models can cause certain heads to focus excessively on a few pixel pairs, particularly in the residual post-normalization setting. To mitigate this, SwinV2 introduces scaled cosine attention, which replaces dot-product similarity with cosine similarity, normalizing attention scores and improving training stability.
SwinV2 also replaces the relative position bias with a log-spaced continuous position bias. This approach is particularly relevant in hierarchical models like SwinV2, where the patch merging operation, detailed in the following paragraph, changes the token resolution at each stage of the network. The log-spaced continuous position bias provides a flexible encoding of spatial relationships that adapts naturally as the resolution changes, allowing the model to maintain meaningful attention patterns at multiple scales. In practice, this helps the model assign more attention to neighboring regions (where fine details or local patterns often reside) while remaining robust to global variations in the visual structure of different light curves.
Following these operations, the model applies patch merging to create a hierarchical representation. This reduces the spatial resolution of the token map while increasing its embedding dimension, inspired by how CNNs progressively downsample feature maps while enriching feature abstraction. During patch merging, each group of $2\times 2$ neighboring tokens is concatenated to form a single new token, halving the height and width of the token map and multiplying the embedding dimension by four. A linear layer then projects these concatenated features into a lower-dimensional space, typically doubling the previous embedding size (e.g., from $64\times 64\times 96$ to $32\times 32\times 192$). By repeating this process across network stages, the model builds a multi-scale or hierarchical representation: early layers retain fine-grained spatial details, while deeper layers encode more abstract and global patterns. This hierarchical structure is designed to enable the model to capture features at different temporal scales, which, in the context of light curves, could correspond to both short-term variations and longer-term trends.
After the hierarchical stages, the final token representations undergo adaptive average pooling, which computes the mean across all tokens of size $\frac{H}{32}\times\frac{W}{32}$, preserving their aggregated embedding information ($8C$). This process allows each image to be represented as a vector of dimension $8C$, which is then passed through a fully connected linear layer for classification or other downstream tasks.
@@ 3.1.1 MACHO labeled
The public Catalog of Variable Stars [^2] MACHO Labeled Data of the Variable Stars. (Alcock et al. 2003) contains labeled light curves from the MACHO survey, which observed variable stars in the Large Magellanic Cloud (LMC) and the Galactic Bulge over multiple years, capturing their photometric variability in the B and R bands. These light curves are represented in magnitude space. The catalog labels were assigned through automated classification procedures based on statistical properties of the light curve morphology, such as period, amplitude, and Fourier decomposition coefficients. While visual verification was commonly employed in some of the MACHO team’s publications (Alcock et al. 1996, 1997a, 1997b, 1999, 2002), there is no evidence in the literature confirming that a systematic visual inspection was performed for the full catalog used in this work. Therefore, we assume that the classification relied solely on the extracted features. It is important to note that we do not use any auxiliary features or pre-computed variables at any stage of the preprocessing or training pipeline, which could otherwise lead to a trivial classification task. This dataset also has a median cadence of $0.75$ days with a standard deviation of $5.03$ days, considering both bands. The labeled portion we use corresponds to the same one used in Donoso-Oliva et al. (2023), which was categorized into six distinct classes, comprising 20 89420\,894 variable stars. We will refer to this as the Full dataset.
Following the approach in Donoso-Oliva et al. (2023), we use scenarios with few samples per class (spc) to train the model and assess its ability to learn from limited labeled data while keeping the test set unchanged. From the Full dataset, we extracted 20 spc and 500 spc as our new training datasets. As in the referenced study, the IDs for each fold were selected randomly from the Full dataset. For example, the IDs for 20 and 500 spc in the first fold differ from those in the second fold. Our training, validation, and test sets in each fold use the same IDs that were used in the referenced study. The test set contains $100$ light curves per class and remains fixed across all scenarios (20 spc, 500 spc, and Full dataset). The training dataset is divided into three stratified folds, maintaining an 80/20 split for training and validation. This results in $96$ and $24$ light curves for the 20 spc scenario, 2 4002\,400 and $600$ for the 500 spc scenario, and 16 23216\,232 and 4 0624\,062 for the Full dataset. Figure 3 illustrates the class distribution in the training and validation sets for the Full dataset.
@@ 3.1.2 ELAsTiCC Challenge
The second dataset used in this study comes from the ELAsTiCC [^3] ELAsTiCC Challenge Data Portal. challenge (Narayan & Team 2023), designed to simulate the observations of the future LSST at the Vera C. Rubin Observatory. It contains 1 845 1461\,845\,146 simulated light curves, each represented across six bands: u, g, r, i, z, and Y, covering the visible and near-infrared spectrum. The data includes multiple observations of each object spread over several simulated years. ELAsTiCC spans a diverse array of transient and variable stellar events, categorized into $32$ distinct classes, with a median cadence of $0.93$ days and a standard deviation of $16.18$ days, considering all six bands. Specifically, we used the first ELAsTiCC campaign, which was used in Cabrera-Vives et al. (2024), and adopted the same class distribution as their study, encompassing $20$ different categories. Additionally, similar to the approach outlined in the referenced study, we used the light curves in flux space and applied the PHOTFLAG [^4] Description of the data variables from the first round of ELAsTiCC. key to filter out saturated data, which also helps exclude spurious detections outside the main event, particularly for transient objects. For clarity, we note that no derived features from the light curves or metadata were used at any stage of the training process. For each light curve, we also considered both alerts and forced photometry points, beginning $30$ days prior to the first alert up to the final detection. Non-detections following the last detection were omitted from the dataset.
Our training, validation, and test sets were the same as those used in Cabrera-Vives et al. (2024). The test set contains 1 0001\,000 light curves from each class, ensuring a balanced representation. The remaining data is divided into five stratified folds, maintaining an 80/20 ratio for training and validation. This results in 1 460 1171\,460\,117 and 365 029365\,029 light curves in the training and validation sets, respectively. Figure 4 shows the distribution of the classes in the training and validation sets.